OpenStudy (anonymous):

1/5 * -10 Multiple , show steps, simlest form start by 1/5 * -10/1 ?

OpenStudy (lalaly):

\[\frac{1}{5} \times \frac{10}{-1} = \frac{1 \times 10}{-1 \times 5} = \frac{10}{-5}=-2\]

OpenStudy (anonymous):

thank you